# Cosmic twins are likely rare: exploring pathways to hadron-quark stars with the same mass as neutron stars

### v1

@@ III Initial data and evolution Methods
In the $M-\rho_{c}$ plane of Fig. 1, the unstable WD configuration B will collapse at constant rest mass, and following cooling, it must ultimately migrate towards one of the two stable configurations with the same rest-mass, i.e., at points C or E. During its evolution, the star moves through configurations that cannot be accurately described by cold equilibrium models, so a priori, it is unclear which of these stable states the star will end up to. To determine the ultimate fate of the unstable WD, we perform multiple simulations of gravitational collapse by perturbing configuration B. As the initial configuration contracts, it heats adiabatically, and via shocks, if the core bounces. Therefore, to end up on a cold degenerate configuration, the star has to cool. We adopt a covariant local effective cooling approach, the details of which we present in App. B. Our cooling model is that of [124], and is characterized by a single parameter–the cooling timescale $\tau_{c}$.
Except for the code that builds our tabulated EOS in Python, and our TOV solver, our computational framework is based on the Einstein Toolkit [125, 126]Cactus [127] and Simfactory [128], employing Carpet for mesh refinement [129]. Post-processing of the simulation data is performed using the kuibit Python package [130].
@@ III.1 Initial data
To vary the degree of violence of the collapse of our initial configuration we consider a set of different initial perturbations, that involve non-zero velocity and/or pressure perturbations. We also consider mass depletion perturbations. We describe these below. After we impose the initial perturbations we do not resolve the Einstein constraints, but we check in all our cases that the constraint violations are always small during the evolution.
@@ III.1.2 Velocity perturbation
where $\vec{v}$ is the coordinate 3-velocity, $\hat{r}$ the radial unit vector, $r$ is the coordinate radius, $R$ the coordinate radius of the star, and $\xi_{v}$ determines the amplitude of the velocity perturbation, and could be either negative (collapsing star) or positive (expanding star). The exponent $\kappa$ controls the radial profile of the perturbation. The case $\kappa=0$ is a particular model with a velocity perturbation independent of position. The choice of a power-law perturbation model is made for simplicity and is not unique. We choose $\kappa\neq 0$ so the perturbation goes smoothly to zero close to the center of the star. This is because for $\xi_{v}=\mathcal{O}(-0.1)$ and $\kappa=0$ the star bounces almost immediately and explodes.
@@ III.1.3 Mass perturbation
Here, $\xi_{m}={\delta M_{0}}/{M_{0}}$ specifies the fraction of rest mass removed or added from/to the original star, and hence, takes a negative value in the case of mass loss. In practice, this is done by defining a threshold rest-mass density below which the rest-mass density of the stars takes the value of the tenuous atmospheric density until the desired amount of mass $\delta M_{0}$ is depleted.
@@
Table 2: The set of simulations conducted in this study using EOS I is detailed below. For each case, the corresponding configuration in the $M-\rho_{c}$ plot for the initial data are listed, along with the perturbation parameters. The last column describes how cooling is performed during the simulation, if enabled. Further details are provided in Sec. IV. Configuration F represents a stable HS with $M_{0}=1.49M_{\odot}$, and is discussed in Sec. IV.3. The simulations based on the last two rows are presented in App. D.
@@ IV.1 White dwarf collapse
Figure 5: Evolution of the maximum density for an unstable WD under 1% pressure depletion. Here, $t_{\rm dyn}=26.46\,\text{ms}$ is the dynamical timescale of the initial WD. The orange and blue dashed lines denote the rest-mass density of the stable compact configurations C and E of Fig. 1, respectively. The black dashed (green) curve illustrates the evolutionary path in the SI (SII) cooling strategy. The inset zooms in on the bounce.
@@
We also conducted a simulation with $\tau_{c}=1.5\,t_{dyn}$ as opposed to $\tau_{c}=3\,t_{dyn}$, to check if the cooling timescale can affect the results. We confirmed that faster cooling only accelerates the collapse, while leaving the final outcome unaffected. Thus, we conclude that the final product of the collapse of an unstable WD is a stable NS with the same total rest mass; this initial data does not result in a stable HS, regardless of the cooling approaches tested.
Figure 6: Same as Fig. 5, but for the case with 90 percent of pressure depletion in the initial data. The line styles and colors are coded as in Fig. 5, and similarly, $t_{\rm dyn}$ is the dynamical timescale of the initial WD
@@ IV.2 Strongly perturbed initial data
Figure 7: Maximum density vs time for the simulations of WD implosion with initial perturbations in both velocity and pressure. Time is scaled with the dynamical timescale of the initial WD. The colors maroon, green, and magenta designate the three values of $\kappa=\{0.5,1,5\}$, respectively. The velocity perturbation is fixed at $\xi_{v}=-0.1$. Dotted lines correspond to $\xi_{p}=-0.01$, while solid lines to $\xi_{p}=-0.90$. All six simulations employ the SII cooling strategy. The dashed lines designate the central density of the stable NS/HS as in Fig. 5.
@@ IV.2.2 Velocity perturbation
Given that the binding energy of the TS is larger than that of the NS with the same rest-mass, it is possible that we must inject energy into the system to overcome the bounce. One way to induce a more violent implosion is by introducing an inward velocity perturbation. This perturbation is modeled with Eq. (7), which consists of two free parameters $\xi_{v}$ and $\kappa$ to be specified. We fix $\xi_{v}=-0.1$, which corresponds to an inward velocity at the surface at 10% of the speed of light, and perform simulations with three different values of $\kappa$, namely $\kappa=\{0.5,1,5\}$. Fig. 7 shows the result of these simulations in the evolution plot of the maximum density, represented by maroon, green, and magenta colors, respectively.
To explore a more diverse range of initial conditions, each of these simulations is performed in two cases: $\xi_{p}=-0.01$ (dotted curves) and $\xi_{p}=-0.90$ (solid curves). In all these simulations, cooling is activated from the beginning, that is, we employ the SII cooling strategy. All these cases yield the same final result as the previous simulation of the collapsing WD; the original bounce near the last stages of the collapse is still unavoidable, and the remaining core is still a stable NS, regardless of the value of $\kappa$. Different values of $\kappa$ can only change the timescale of the collapse. Among these three choices, $\kappa=0.5$ and $\kappa=5$ result in the fastest and the slowest processes, respectively. Moreover, the combination of the velocity perturbation with a larger pressure depletion in the initial data leads to a faster implosion in all cases.
As the last set of simulations for the collapse of an unstable WD, we also explore velocity perturbations with larger values of $\xi_{v}$. Fig. 8 presents these simulations with $\xi_{v}=-0.3$ (dotted curve) and $\xi_{v}=-0.4$ (solid curve). Given that all choices of $\kappa$ led to the same ultimate result, we only run these simulations for the case of $\kappa=1$. Both simulations have 90 percent of pressure depletion, $\xi_{p}=-0.90$, and are conducted in the SII. These represent the most extreme initial conditions examined in this study. However, the result is once again the stable NS at configuration C. The bounce still manifests after the exponential collapse, although the process is even faster than in the previous cases. The general outcome of these simulations suggests that even extreme pressure and velocity perturbations cannot change the eventual fate of the collapsing WD. The outcome is always a NS.
@@ IV.3 Mass loss
In this section, we investigate the scenario where the NS maximum mass is somewhat lower than the Chandrasekhar limit so that a stellar core or a WD at the Chandrasekhar limit collapses to the stable hybrid star branch above the twin star mass regime. If this initial hybrid star experiences some mass loss, e.g., due to strong winds, it could settle into a TS instead of an NS.
To model this scenario, we take the hybrid star configuration F in Fig. 9, and deplete a small fraction of its mass as described in Eq. (8). The initial rest mass of this configuration is $M_{0}=1.49\,M_{\odot}$, and exceeds $M_{ch}$. The initial central rest-mass density is $\rho_{c}=1.63\times 10^{15}\,\textrm{g}\,\textrm{cm}^{-3}$ and the radius is $R=11.02\,\textrm{km}$. By setting $\xi_{m}=-0.033$, the rest mass of the star becomes $M_{0}\approx M_{0,s}$, aligning with the rest masses of stable stars A, C, and E of Fig. 1. This initial perturbation is fairly strong compared to a slow loss of mass through winds, but it allows us to get a glimpse into whether this pathway to forming TSs is viable. We checked the constraint violations after the mass loss remain small throughout the simulations we performed (see discussion in App. E).
Figure 9: A zoomed-in view of the NS/HS part of the Fig. 1. Configuration F is a stable HS with a mass higher than the other three configurations indicated. The green shaded area roughly denotes the mass range for TSs. The red (blue) curve corresponds the ADM (rest) mass.
@@
Figure 11: Rest mass as a function of central density for a sequence of TOV stars with EOS I (dotted) and EOS II (solid). The dashed line represents the Chandrasekhar mass. Notably, there is no TS in EOS II with a rest mass close to $M_{ch}$. In the density range relevant to WDs, the two curves exhibit almost complete overlap. Thus, initial unstable WD configuration (B) shares similar characteristics in both EOSs.
Fig. 10 illustrates the results. Four simulations are shown in this figure; the stable configurations C, E, and F (without perturbation) are displayed by the orange, blue, and violet solid lines, respectively. The black solid curve represents the evolution of the maximum density of a star beginning at configuration F but perturbed by $\xi_{m}=-0.033$. Interestingly, this plot demonstrates that the last scenario successfully ends up with the formation of a stable HS in the twin star regime. As a result of this large initial perturbation, two types of oscillations can be seen around the maximum density of the stable HS shown in blue. The first type exhibits oscillations with a period of $\sim t_{dyn}$, which gradually disappears after a few dynamical timescales. The second type of oscillation has a larger domain with ±$\sim\pm 10\,\%$ of deviations around the mean value of the central rest-mass density of the TS E. The latter type of oscillation has been seen before in simulations with EOSs that exhibit a strong phase transition, and is most likely due to a significant portion of the mass of the star transitioning in and out of the density range in the hadron-to-quark phase transition, where the stiffness of the EOS changes rapidly leading to collapse (EOS softens) and bounce (EOS stiffens). We expect that a gradual loss of rest-mass would make such oscillations disappear, but the main result of the final configuration being a TS with the same rest mass should hold. The important finding here is that the maximum central density remains above the phase transition region as shown in Fig. 10, and its mean is that of the TS with the same rest mass.
@@ IV.4 Hybrid star formation
The simulation we perform here is the only one in this work based on EOS II listed in Table 1. Fig. 11 compares the $M_{0}-\rho_{c}$ curves resulting from these two EOSs. The TOV sequence resulting from EOS II has an NS peak lower than $M_{ch}$. Therefore, there are no TSs with rest mass $M_{0,s}$.
@@
Figure 12: The evolution of the maximum density of a star beginning as an unstable WD with a rest mass $M_{0}=M_{0,s}$. The blue line represents the central density of the stable HS with this rest mass, and the phase transition range of densities is indicated by the shaded gray area. Here $t_{\rm dyn}$ represents dynamical timescale of the initial WD. The inset zooms in on the final stages of collapse, indicating the convergence of maximum density to the central density of the stable HS.
@@ V Conclusions
In this work, we explored the formation of TSs by conducting hydrodynamical simulations in full general relativity. We constructed piecewise polytropic EOSs that approximate realistic EOSs with a hadron-to-quark phase transition, and we extended them down to the EOS of a white dwarf. Using our EOSs, we performed multiple simulations of the collapse of an unstable WD under different initial conditions with varying degrees of extremality of initial perturbations in pressure and/or velocity. Given that total rest mass is conserved, the ultimate product of the collapse could either be a stable NS or a stable HS. The general finding from all our simulations is that the unstable WD always collapses into a stable NS, regardless of how extreme the initial conditions are. This overall result remains the same even under extreme initial perturbations.
@@
If TSs exist, there should be at least one pathway for them to form. Following the standard theory of NS formation, a natural path could be the formation of a more massive HS that subsequently loses a small amount of mass in the form of winds. Another potential avenue for mass loss would be a “grazing” collision of a massive HS with a black hole where the massive HS undergoes an episode of mass loss and then flies away. Such a scenario might take place in a dense stellar cluster. Our simulations show that forming HSs heavier than in the twin star mass regime is possible, and that forming a TS through a heavier HS experiencing mass loss is a viable path. However, even this scenario appears that it would require some fine-tuning for TSs to form, which would place strong limits on the abundance of TSs. The fine tuning for the formation of twin stars results because of the limited range of masses over which they are predicted to reside by existing EOSs with hadron-to-quark phase transitions, and hence they should be fairly rare objects. This conclusion and the challenge in producing TSs as demonstrated by our simulations suggest that if a HS star was involved in GW170817, then it likely was not a twin star.
@@
We point out that our work is idealized in several ways and has a number of caveats. First, we do not treat realistic neutrino effects or magnetic fields, and the perturbations studied here represent a simplified version of processes that may occur inside the dense core of massive stars during the last stages of their evolution. In particular, scenarios involving mass loss due to winds and mass gain resulting from the infall of matter from outer shells of the initial progenitor are expected to happen in nature. The results of our simulations indicate that these perturbations play a notable role in the potential formation of stable TSs, assuming that they exist in nature. Thus, it is important to treat these self-consistently. Nevertheless, our conclusion that NSs tend to be the preferred outcome in the twin star mass range should hold, not only because this is what our simulations demonstrate, but also because of the very small range of masses over which twin stars are predicted to exist by current EOSs, which in turn requires a delicate balance of mass loss and mass gain.
@@
Another potential pathway to forming HSs involves mergers of binary NSs or WD-NS [140, 124]. Exploring all these scenarios and including more realistic physics is left for future work. Most work in the area of TSs revolves around either observational constraints on or properties of the EOS. Our work here points out that there is another important problem surrounding TSs: their formation and abundance in nature.

### v2

@@ III Initial data and evolution Methods
In the $M-\rho_{c}$ plane of Fig. 1, the unstable WD configuration B will collapse at constant rest mass, and following cooling, it must ultimately migrate toward one of the two stable configurations with the same rest mass, i.e., at points C or E. During its evolution, the star moves through configurations that cannot be accurately described by cold equilibrium models, so a priori, it is unclear which of these stable states the star will end up to. To determine the ultimate fate of the unstable WD, we perform multiple simulations of gravitational collapse by perturbing configuration B. As the initial configuration contracts, it heats adiabatically, and via shocks, if the core bounces. Therefore, to end up on a cold degenerate configuration, the star has to cool. We adopt a covariant local effective cooling approach, the details of which we present in App. B. Our cooling model is that of [124], and is characterized by a single parameter–the cooling timescale $\tau_{c}$.
Except for the code that builds our tabulated EOS in Python, and our TOV solver, our computational framework is based on the Einstein Toolkit [125, 126]Cactus [127] and Simfactory [128], employing Carpet for mesh refinement [129]. Postprocessing of the simulation data is performed using the kuibit Python package [130].
@@ III.1 Initial data
To vary the degree of violence of the collapse of our initial configuration we consider a set of different initial perturbations that involve nonzero velocity and/or pressure perturbations. We also consider mass depletion perturbations. We describe these below. After we impose the initial perturbations we do not resolve the Einstein constraints, but we check in all our cases that the constraint violations are always small during the evolution.
@@ III.1.2 Velocity perturbation
where $\vec{v}$ is the coordinate 3-velocity, $\hat{r}$ is the radial unit vector, $r$ is the coordinate radius, $R$ is the coordinate radius of the star, and $\xi_{v}$ determines the amplitude of the velocity perturbation and could be either negative (collapsing star) or positive (expanding star). The exponent $\kappa$ controls the radial profile of the perturbation. The case $\kappa=0$ is a particular model with a velocity perturbation independent of position. The choice of a power-law perturbation model is made for simplicity and is not unique. We choose $\kappa\neq 0$ so the perturbation goes smoothly to zero close to the center of the star. This is because for $\xi_{v}=\mathcal{O}(-0.1)$ and $\kappa=0$ the star bounces almost immediately and explodes.
@@ III.1.3 Mass perturbation
Here, $\xi_{m}={\delta M_{0}}/{M_{0}}$ specifies the fraction of rest mass removed from or added to the original star and, hence, takes a negative value in the case of mass loss. In practice, this is done by defining a threshold rest-mass density below which the rest-mass density of the stars takes the value of the tenuous atmospheric density until the desired amount of mass $\delta M_{0}$ is depleted.
@@
Table 2: The set of simulations conducted in this study using EOS I is detailed below. For each case, the corresponding configuration in the $M-\rho_{c}$ plot for the initial data is listed, along with the perturbation parameters. The last column describes how cooling is performed during the simulation, if enabled. Further details are provided in Sec. IV. Configuration F represents a stable HS with $M_{0}=1.49M_{\odot}$, and is discussed in Sec. IV.3. The simulations based on the last two rows are presented in App. D.
@@ IV.1 White dwarf collapse
Figure 5: Evolution of the maximum density for an unstable WD under 1% pressure depletion. Here, $t_{\rm dyn}=26.46\,\text{ms}$ is the dynamical timescale of the initial WD. The orange and blue dashed lines denote the rest-mass density of the stable compact configurations C and E of Fig. 1, respectively. The black dashed (green) curve illustrates the evolutionary path in the SI (SII) cooling strategy. The inset enlarges the bounce.
@@
We also conducted a simulation with $\tau_{c}=1.5\,t_{dyn}$ as opposed to $\tau_{c}=3\,t_{dyn}$, to check if the cooling timescale can affect the results. We confirmed that faster cooling only accelerates the collapse, while leaving the final outcome unaffected. Thus, we conclude that the final product of the collapse of an unstable WD is a stable NS with the same total rest mass; these initial data do not result in a stable HS, regardless of the cooling approaches tested.
Figure 6: The same as Fig. 5, but for the case with 90 percent of pressure depletion in the initial data. The line styles and colors are coded as in Fig. 5, and similarly, $t_{\rm dyn}$ is the dynamical timescale of the initial WD
@@ IV.2 Strongly perturbed initial data
Figure 7: Maximum density versus time for the simulations of WD implosion with initial perturbations in both velocity and pressure. Time is scaled with the dynamical timescale of the initial WD. The colors maroon, green, and magenta designate the three values of $\kappa=\{0.5,1,5\}$, respectively. The velocity perturbation is fixed at $\xi_{v}=-0.1$. Dotted lines correspond to $\xi_{p}=-0.01$, while solid lines to $\xi_{p}=-0.90$. All six simulations employ the SII cooling strategy. The dashed lines designate the central density of the stable NS/HS as in Fig. 5.
@@ IV.2.2 Velocity perturbation
Given that the binding energy of the TS is larger than that of the NS with the same rest mass, it is possible that we must inject energy into the system to overcome the bounce. One way to induce a more violent implosion is by introducing an inward velocity perturbation. This perturbation is modeled with Eq. (7), which consists of two free parameters $\xi_{v}$ and $\kappa$ to be specified. We fix $\xi_{v}=-0.1$, which corresponds to an inward velocity at the surface at 10% of the speed of light, and perform simulations with three different values of $\kappa$, namely $\kappa=\{0.5,1,5\}$. Fig. 7 shows the result of these simulations in the evolution plot of the maximum density, represented by maroon, green, and magenta colors, respectively.
To explore a more diverse range of initial conditions, each of these simulations is performed in two cases: $\xi_{p}=-0.01$ (dotted curves) and $\xi_{p}=-0.90$ (solid curves). In all these simulations, cooling is activated from the beginning, that is, we employ the SII cooling strategy. All these cases yield the same final result as the previous simulation of the collapsing WD; the original bounce near the last stages of the collapse is still unavoidable, and the remaining core is still a stable NS, regardless of the value of $\kappa$. Different values of $\kappa$ can change only the timescale of the collapse. Among these three choices, $\kappa=0.5$ and $\kappa=5$ result in the fastest and the slowest processes, respectively. Moreover, the combination of the velocity perturbation with a larger pressure depletion in the initial data leads to a faster implosion in all cases.
As the last set of simulations for the collapse of an unstable WD, we also explore velocity perturbations with larger values of $\xi_{v}$. Fig. 8 presents these simulations with $\xi_{v}=-0.3$ (dotted curve) and $\xi_{v}=-0.4$ (solid curve). Given that all choices of $\kappa$ led to the same ultimate result, we run these simulations only for the case of $\kappa=1$. Both simulations have 90 percent of pressure depletion, $\xi_{p}=-0.90$, and are conducted in the SII. These represent the most extreme initial conditions examined in this study. However, the result is once again the stable NS at configuration C. The bounce still manifests after the exponential collapse, although the process is even faster than in the previous cases. The general outcome of these simulations suggests that even extreme pressure and velocity perturbations cannot change the eventual fate of the collapsing WD. The outcome is always a NS.
@@ IV.3 Mass loss
In this section, we investigate the scenario where the NS maximum mass is somewhat lower than the Chandrasekhar limit so that a stellar core or a WD at the Chandrasekhar limit collapses to the stable hybrid star branch above the twin star mass regime. If this initial hybrid star experiences some mass loss, e.g., due to strong winds, it could settle into a TS instead of a NS.
To model this scenario, we take the hybrid star configuration F in Fig. 9, and deplete a small fraction of its mass as described in Eq. (8). The initial rest mass of this configuration is $M_{0}=1.49\,M_{\odot}$, and exceeds $M_{ch}$. The initial central rest-mass density is $\rho_{c}=1.63\times 10^{15}\,\textrm{g}\,\textrm{cm}^{-3}$ and the radius is $R=11.02\,\textrm{km}$. By setting $\xi_{m}=-0.033$, the rest mass of the star becomes $M_{0}\approx M_{0,s}$, aligning with the rest masses of stable stars A, C, and E in Fig. 1. This initial perturbation is fairly strong compared to a slow loss of mass through winds, but it allows us to get a glimpse into whether this pathway to forming TSs is viable. We checked the constraint violations after the mass loss remain small throughout the simulations we performed (see discussion in App. E).
Figure 9: An enlarged view of the NS/HS part of Fig. 1. Configuration F is a stable HS with a mass higher than the other three configurations indicated. The green shaded area roughly denotes the mass range for TSs. The red (blue) curve corresponds the ADM (rest) mass.
@@
Figure 11: Rest mass as a function of central density for a sequence of TOV stars with EOS I (dotted line) and EOS II (solid line). The dashed line represents the Chandrasekhar mass. Notably, there is no TS in EOS II with a rest mass close to $M_{ch}$. In the density range relevant to WDs, the two curves exhibit almost complete overlap. Thus, the initial unstable WD configuration (B) shares similar characteristics in both EOSs.
Fig. 10 illustrates the results. Four simulations are shown in this figure; the stable configurations C, E, and F (without perturbation) are displayed by the orange, blue, and violet solid lines, respectively. The black solid curve represents the evolution of the maximum density of a star beginning at configuration F but perturbed by $\xi_{m}=-0.033$. Interestingly, this plot demonstrates that the last scenario successfully ends up with the formation of a stable HS in the twin star regime. As a result of this large initial perturbation, two types of oscillations can be seen around the maximum density of the stable HS shown in blue. The first type exhibits oscillations with a period of $\sim t_{dyn}$, which gradually disappears after a few dynamical timescales. The second type of oscillation has a larger domain with ±$\sim\pm 10\,\%$ of deviations around the mean value of the central rest-mass density of the TS E. The latter type of oscillation has been seen before in simulations with EOSs that exhibit a strong phase transition [91, 140], and is likely due to a significant portion of the mass of the star transitioning in and out of the density range in the hadron-to-quark phase transition, where the stiffness of the EOS changes rapidly leading to collapse (EOS softens) and bounce (EOS stiffens). We expect that a gradual loss of rest mass would make such oscillations disappear, but the main result of the final configuration being a TS with the same rest mass should hold. The important finding here is that the maximum central density remains above the phase transition region as shown in Fig. 10, and its mean is that of the TS with the same rest mass.
@@ IV.4 Hybrid star formation
The simulation we perform here is the only one in this work based on EOS II listed in Table 1. Fig. 11 compares the $M_{0}-\rho_{c}$ curves resulting from these two EOSs. The TOV sequence resulting from EOS II has a NS peak lower than $M_{ch}$. Therefore, there are no TSs with rest mass $M_{0,s}$.
@@
Figure 12: The evolution of the maximum density of a star beginning as an unstable WD with a rest mass $M_{0}=M_{0,s}$. The blue line represents the central density of the stable HS with this rest mass, and the phase transition range of densities is indicated by the shaded gray area. Here $t_{\rm dyn}$ represents dynamical timescale of the initial WD. The inset enlarges the final stages of collapse, indicating the convergence of maximum density to the central density of the stable HS.
@@ V Conclusions
In this work, we explored the formation of TSs by conducting hydrodynamical simulations in full general relativity. We constructed piecewise polytropic EOSs that approximate realistic EOSs with a hadron-to-quark phase transition, and we extended them down to the EOS of a white dwarf. Using our EOSs, we performed multiple simulations of the collapse of an unstable WD under different initial conditions with varying degrees of extremality of initial perturbations in pressure and/or velocity. Given that total rest mass is conserved, the ultimate product of the collapse could be either a stable NS or a stable HS. The general finding from all our simulations is that the unstable WD always collapses into a stable NS, regardless of how extreme the initial conditions are. This overall result remains the same even under extreme initial perturbations.
@@
If TSs exist, there should be at least one pathway for them to form. Following the standard theory of NS formation, a natural path could be the formation of a more massive HS that subsequently loses a small amount of mass in the form of winds. Another potential avenue for mass loss would be a “grazing” collision of a massive HS with a black hole where the massive HS undergoes an episode of mass loss and then flies away. Such a scenario might take place in a dense stellar cluster. Our simulations show that forming HSs heavier than in the twin star mass regime is possible, and that forming a TS through a heavier HS experiencing mass loss is a viable path. However, even this scenario appears that it would require some fine-tuning for TSs to form, which would place strong limits on the abundance of TSs. The fine-tuning for the formation of twin stars results because of the limited range of masses over which they are predicted to reside by existing EOSs with hadron-to-quark phase transitions, and, hence, they should be fairly rare objects. This conclusion and the challenge in producing TSs as demonstrated by our simulations suggest that if a HS star was involved in GW170817, then it likely was not a twin star.
@@
We point out that our work is idealized in several ways and has a number of caveats. First, we do not treat realistic neutrino effects or magnetic fields, and the perturbations studied here represent a simplified version of processes that may occur inside the dense core of massive stars during the last stages of their evolution. In particular, scenarios involving mass loss due to winds and mass gain resulting from the infall of matter from outer shells of the initial progenitor are expected to happen in nature. The results of our simulations indicate that these perturbations play a notable role in the potential formation of stable TSs, assuming that they exist in nature. Thus, it is important to treat these self-consistently. Nevertheless, our conclusion that NSs tend to be the preferred outcome in the twin star mass range should hold, not only because this is what our simulations demonstrate, but also because of the very small range of masses over which twin stars are predicted to exist by current EOSs, which, in turn, requires a delicate balance of mass loss and mass gain.
@@
The effects of rotation was not treated here, but such effects on HSs have been studied [55, 141, 142, 71]. The centrifugal barrier due to rotation is likely to lower the density compared to a nonrotating configuration for a given rest mass, thereby making matter stay well below the hadron-to-quark phase transition density. In addition, for some hybrid EOSs rotation has the effect of increasing the neutron star maximum mass higher than that of the third family [36]. However, these results hold for cold configurations.
Temperature is another quantity that could affect stellar properties [143]; notably, phase transitions may occur at finite rather than zero temperature [144]. The EOS we adopted in this work exhibits a hadron-to-quark phase transition at zero temperature. One scenario that can potentially change the results is the possibility of an EOS with a phase transition at finite temperatures [95]. In this scenario, phase transition occurs at a lower density when the star is still hot [145]. Therefore, one can also think about the formation of TSs considering such a EOS with a finite-temperature phase transition. We will explore this possibility in future work.
Another potential pathway to forming HSs involves mergers of binary NSs or WD-NS [146, 124, 147]. Exploring all these scenarios and including more realistic physics is left for future work. Most work in the area of TSs revolves around either observational constraints on or properties of the EOS. Our work here points out that there is another important problem surrounding TSs: their formation and abundance in nature.